# TextMine: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action

### v2

@@ Abstract
Humanitarian Mine Action (HMA) addresses the challenge of detecting and removing landmines from conflict regions. Much of the life-saving operational knowledge produced by HMA agencies is buried in unstructured reports, limiting the transferability of information between agencies. To address this issue, we propose TextMine: the first dataset, evaluation framework and ontology-guided large language model (LLM) pipeline for knowledge extraction in the HMA domain. TextMine structures HMA reports into (subject, relation, object)-triples, thus creating domain-specific knowledge. To ensure real-world relevance, we created the dataset in collaboration with Cambodian Mine Action Center (CMAC). We further introduce a bias-aware evaluation framework that combines human-annotated triples with an LLM-as-Judge protocol to mitigate position bias in reference-free scoring. Our experiments show that ontology-aligned prompts improve extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format adherence by 20.9% compared to baseline models. We publicly release the dataset and code. [^1] https://anonymous.4open.science/r/TextMine_Code_Dataset-D974
TextMine: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action
Chenyue Zhou1,2   Gürkan Solmaz1   Flavio Cirillo1   Kiril Gashteovski1,4   Jonathan Fürst3 1NEC Laboratories Europe, Germany, 2University of Stuttgart, Germany, 3Zurich University of Applied Sciences, Switzerland 4CAIR, Ss. Cyril and Methodius University of Skopje, North Macedonia st185961@stud.uni-stuttgart.de guerkan.solmaz,flavio.cirillo,kiril.gashteovski@neclab.eu jonathan.fuerst@zhaw.ch
@@ 1 Introduction
In this work, we propose TextMine: the first dataset, evaluation framework and pipeline for IE for the HMA domain. The task is to extract information from HMA reports in the form of (subject, relation, object)-triples, thus creating structured knowledge for the HMA domain (Fig. 1). Prior data and benchmarks (Mihindukulasooriya et al., 2023, inter alia) rely on toy data schemas (i.e., ontologies), which limits their applicability to the complex and specialized HMA documents. By contrast, we created the data in collaboration with the Cambodian Mine Action Center (CMAC). CMAC has been initiated by the UN more than 30 years ago, is one of the world’s most experienced and effective demining organizations. Its collaborations within the UN and the Geneva International Centre for Humanitarian Demining (GICHD), ensure that the data is useful in real-world HMA applications beyond Cambodia Landmine and Cluster Munition Monitor (2023); United Nations Development Programme in Cambodia (2021). While there is evidence that ontology-guided methods are effective (Cauter and Yakovets, 2024), they are limited to single-sentence inputs and overlook context-level reasoning. TextMine addresses this gap by enabling context-level reasoning, while aggregating a set of HMA-related ontologies. Finally, no framework is bias-aware, which might distort the evaluation results, a gap that TextMine also addresses.
To sum up, our contributions are: 1 Data: we introduce a curated dataset, curated ontology and annotated ground-truth data for humanitarian demining operations, systematically categorizing operational entities and relationships. 2 Evaluation: We evaluate extracted triples against our annotated dataset and introduce a bias-aware LLM-as-Judge framework for reference-free scoring. Experiments on closed and open LLMs show that position bias skews rankings. 3 Knowledge Extraction: We propose a prompt-based pipeline that combines layout-aware document chunking, ontology-guided extraction, and multi-perspective evaluation. To our knowledge, this is the first LLM application of knowledge extraction for HMA. 4 In-Context Learning Optimization: We found that prompts enriched with ontology-aligned examples improve triple extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format conformance by up to 20.9% compared to baseline LLMs. These findings provide practical insights for prompt construction.
@@ 2 Benchmark Dataset Creation
We illustrate the process of creating the annotated humanitarian dataset in Fig. 2. We curated 120 online available technical reports: 60 reports from the Cambodian Mine Action Center (CMAC) and 60 from Geneva International Centre for Humanitarian Demining (GICHD) websites 1. The reports are all in the PDF format. The curation is based on relevance to the humanitarian demining, language (English), and recency. From the initial humanitarian dataset of 120 PDF documents 1, we filtered a dataset of five most recent mine action reports from CMAC of overall 233 pages (see Table 1) 2.
Table 1: Dataset Statistics for TextMine Corpus
@@
Simultaneously, we selected domain-specific ontologies through working with domain experts, by first incorporating data models from the Information Management System for Mine Action (IMSMA) and then manually filtering them for high quality 3. We performed a survey of demining related ontologies together with domain experts and we incorporate six ontologies from Information Management System for Mine Action (IMSMA) Core [^2] https://www.gichd.org/our-response/information-management/imsma-core/. These ontologies are used for demining information system but not strictly for HMA, We then manually filter out concepts and relations unrelated to HMA. Furthermore, we add a more general humanitarian domain ontology from Empathi (Gaur et al., 2019) to make the overall HMA ontology more comprehensive. As a result, HMA ontology integrates seven ontologies (160 entity types, 86 relation types) covering diverse aspects of HMA (see details in Appendix B)..
The annotation prompt generation 4 first parses reports into text chunks, then systematically combines each chunk with ontology templates. Each template specifies the relation types for a specific domain (e.g., mine action events, land contamination, mine clearance). This generates 2,520 prompts (360 chunks $\times$ 7 templates), where each prompt contains both report context and ontology specifications for that domain. We randomly select 100 prompts 5 out of the 2,520, and apply them with the same prompts using GPT-4o and Llama3-70B 6 to initially annotate the data as the “LLM annotations 7.
For each prompt and data chunk pair, each LLM generates a set of knowledge triples as outputs. Then, a human annotator reviewed these outputs, filtering out incorrect knowledge triples and aggregating the valid ones in a new set. Incorrect knowledge triples can include triples following wrong format, mixed order of entities, hallucinated entities/relations (e.g., triples that are not included in the data chunk) and so on. In addition, there exist duplicate triples which are removed to create the final reference set. Those included 1,095 unique triples across the 100 prompts 8. The end result of the human annotation is a set of knowledge triples which are relevant and correct without hallucination or formatting problems, faithful to the given data chunk. This clean set comprises the “annotated test data” 9. The annotated test data can serve as the ground-truth for evaluations of performance of different techniques and LLMs.
@@ 3.1 Humanitarian Mine Action Task
Figure 3: TextMine Overview. Reports are preprocessed into paragraph chunks, used as test inputs during inference. Each chunk is combined with an instruction template and ontology, then passed through LLMs for triple extraction. We apply a multi-perspective evaluation using both reference-based and reference-free methods. Extracted triples are stored in a database, queried by developers and HMA domain experts.
@@ 3.3 Multi-Perspective Evaluation
Triple extraction accuracy serves as the primary metric, as it directly measures how well models extract knowledge triples from text. However, accuracy alone does not fully capture model reliability. The hallucination rate evaluates faithfulness by detecting extraneous or fabricated information, while format conformance ensures that outputs adhere to a syntactically valid structure, enabling seamless integration into downstream applications. We employ N-gram matching-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to assess the extracted triples. Additionally, we incorporate BERTScore (Zhang et al., 2019), which leverages word embeddings to capture semantic similarity beyond lexical overlap.
To enhance the accuracy of our metrics, we first apply stemming and lemmatization by NLTK to normalize morphological variations. We then compute the accuracy metrics against the manually annotated test set.
@@
We try to find the optimal LLM Judge setting by conducting systematic ranking experiments and analyzing correlations between the LLMs judged ranking and reference-based rankings. For these ranking experiments, we design Judge Prompts that instruct LLMs on evaluation criteria. We use five models: Mistral-7B Jiang et al. (2023), Llama3-8B Grattafiori et al. (2024), Gemma2-9B Team et al. (2024), LLaMA3-70B and GPT-4o Hurst et al. (2024), as extraction models. We rank five responses from five models using GPT-4o, Llama3.1-70B, and Llama3.3-70B as our judge models. the correlation between different judge models and methods are included in Table 2. The judge prompts follow a fixed template with seven placeholders, where the ontology placeholder represents entity and relation types. Formally, let the input set for the LLM judge prompts be $\{O,C,R_{m_{1}},R_{m_{2}},R_{m_{3}},R_{m_{4}},R_{m_{5}}\}$, where $O$ is the ontology set, $C$ is the set of test contexts, and $R_{m_{1}},\cdots,R_{m_{5}}$ are the five sets of candidate answers from the five extractor models. The judge LLM produces a verdict (output as a ranking):
@@ 4 Experimental Results
We assess the effectiveness of our knowledge triple extraction method through reference-based and reference-free evaluations. Reference-based evaluation compares extracted triples against our curated dataset, while reference-free evaluation relies on LLM judges to assess generated triples without reference data. A key aspect of our analysis is examining the correlation between these two evaluation paradigms to evaluate the reliability of LLM-based judgments. Additionally, we investigate how different prompt strategies influence extraction performance across models.
@@
Figure 4: The combined visualization illustrates the impact of model selection and prompt strategy on extraction performance. Abbriviations on x-axis are about One-shot with: RS = Random Sentences; RP = Random Paragraphs; OS = Ontology-Aligned Sentence; OP = Ontology-Aligned Paragraph. The top Combined Score is achieved by Llama3-70B (93.24) closely followed by GPT-4o (93.13), both with OS prompt setting.
Figure 5 further breaks down the performance of five models across four accuracy evaluation metrics: BLEU, ROUGE, METEOR, and BERTScore. OS demonstration prompts consistently result in the best accuracy across all four metrics and all five models, highlighting their effectiveness for the triple extraction task. BLEU scores peak with OS prompts, with GPT-4o achieving the best performance. ROUGE results show a similar trend, with GPT-4o and Llama3-70B excelling in the OS prompt setting. METEOR follows the same pattern as BLEU and ROUGE, reinforcing the advantages of OS prompts. BERTScore, which measures semantic similarity, shows high clustering across models, suggesting minimal differentiation in performance.
Overall, OS demonstration prompts consistently enhance extraction accuracy across all models and metrics.
@@
5 Related Work
@@ 6 Conclusion
TextMine addresses the need for knowledge extraction from HMA reports by LLMs and domain ontologies to transform unstructured technical reports into structured knowledge triples. This paper introduces the curated data of HMA reports, ontology, and LLM pipeline, aiming to enable innovations for mine action and natural language processing.
@@
While TextMine provides the first dataset, evaluation framework and pipeline that demonstrates the feasibility of an LLM-driven extraction pipeline in the specialized domain of humanitarian demining, several limitations are considered to be addressed in the future, including the further extension of the dataset. Our evaluation relies on 100 prompt–response examples (yielding 1,095 unique triples). Assembling and annotating demining data demands extensive domain expertise, so even this modest set offers valuable proof-of-concept insights, but we plan to extend to larger, multilingual collections in future work. Second, this work is focues on reports that are written only on English. In the future, we plan to extend the work to include other languages, because mine-action data are often produced in multiple languages, such as Arabic, French or Ukrainian, depending on the conflict region. Making TextMine multilingual is, therefore, essential to ensure global applicability and inclusiveness.
@@
Humanitarian demining is a high importance decision making domain where incorrect or hallucinated triples can misinform planning and lead to wasted time and resources. To address this, TextMine combines reference based validation with a bias aware LLM as Judge framework and publishes all prompts and decoding settings for full transparency. We also engage CMAC and ICRC landmine clearance experts throughout development to review and validate outputs. TextMine is not used to locate mines, as safety remains governed by established GICHD standards [^4] GICHD Mine Action Standards, https://www.gichd.org/our-response/mine-action-standards/, and instead supports expert analysis and planning. By documenting our methods and keeping an expert in the loop, we aim to minimize misinformation and ensure responsible AI deployment in demining operations.
The dataset used in this study consists of publicly available or institutionally provided humanitarian demining reports. These reports were reviewed to ensure they do not contain personally identifiable information (PII) or offensive content. Where necessary, documents were anonymized or filtered to remove sensitive information. Our usage of the data adheres to privacy standards and is strictly confined to research contexts.
All datasets used in this study were accessed under conditions permitting research use. The curated demining report dataset we constructed is intended solely for academic and research purposes and complies with the original access and licensing conditions. The ontology and pipeline components developed in TextMine are likewise designed for research and evaluation within humanitarian domains. We do not support or promote deployment of these artifacts in operational or commercial contexts without further validation and ethical review.

### v3

# TextMineX: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action

@@ Abstract
Humanitarian Mine Action (HMA) addresses the challenge of detecting and removing landmines from conflict regions. Much of the life-saving operational knowledge produced by HMA agencies is buried in unstructured reports, limiting the transferability of information between agencies. To address this issue, we propose TextMineX: the first dataset, evaluation framework and ontology-guided large language model (LLM) pipeline for knowledge extraction from text in the HMA domain. TextMineX structures HMA reports into (subject, relation, object)-triples, thus creating domain-specific knowledge. To ensure real-world relevance, we utilized the dataset from our collaborator Cambodian Mine Action Centre (CMAC). We further introduce a bias-aware evaluation framework that combines human-annotated triples with an LLM-as-Judge protocol to mitigate position bias in reference-free scoring. Our experiments show that ontology-aligned prompts improve extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format adherence by 20.9% compared to baseline models. We publicly release the dataset and code [^1] https://github.com/nec-research/TextMineX.
TextMineX: Data, Evaluation Framework and Ontology-guided LLM Pipeline for Humanitarian Mine Action
Chenyue Zhou1,2,3   Gürkan Solmaz1   Flavio Cirillo1   Kiril Gashteovski1,4   Jonathan Fürst3 1NEC Laboratories Europe, Germany 2University of Stuttgart, Germany 3VAGO Solutions, Germany 4Zurich University of Applied Sciences, Switzerland 5CAIR, Ss. Cyril and Methodius University of Skopje, North Macedonia zhou@vago-solutions.ai, jonathan.fuerst@zhaw.ch guerkan.solmaz,flavio.cirillo,kiril.gashteovski@neclab.eu
@@ 1 Introduction
In this work, we propose TextMineX: the first dataset, evaluation framework and pipeline for IE for the HMA domain. The task is to extract information from HMA reports in the form of (subject, relation, object)-triples, thus creating structured knowledge for the HMA domain (Fig. 1). The application provides demining technical information in the form of knowledge triples that can be stored, shared, and queried across demining agencies. Prior data and benchmarks (Mihindukulasooriya et al., 2023, inter alia) rely on toy data schemas (i.e., ontologies), which limits their applicability to the complex and specialized HMA documents. By contrast, we utilized the publicly available technical reports from our collaborator Cambodian Mine Action Centre (CMAC). CMAC has been initiated by the UN more than 30 years ago, is one of the world’s most experienced and effective demining organizations. Its collaborations within the UN and the Geneva International Centre for Humanitarian Demining (GICHD), ensure that the data is useful in real-world HMA applications beyond Cambodia Landmine and Cluster Munition Monitor (2023); United Nations Development Programme in Cambodia (2021). While there is evidence that ontology-guided methods are effective (Cauter and Yakovets, 2024), they are limited to single-sentence inputs and overlook context-level reasoning. TextMineX addresses this gap by enabling context-level reasoning, while aggregating a set of HMA-related ontologies. Finally, no framework is bias-aware, which might distort the evaluation results, a gap that TextMineX also addresses.
To sum up, our contributions are: 1 Data: we introduce a curated dataset, curated ontology and annotated ground-truth data for humanitarian demining operations, systematically categorizing operational entities and relationships. 2 Evaluation: We evaluate extracted triples against our annotated dataset and introduce a bias-aware LLM-as-Judge framework for reference-free scoring. Experiments on closed and open LLMs show that position bias skews rankings.
3
Knowledge Extraction: We propose a prompt-based pipeline that combines layout-aware document chunking, ontology-guided extraction, and multi-perspective evaluation. To our knowledge, this is the first LLM application of knowledge extraction for HMA. 4 In-Context Learning Optimization: We found that prompts enriched with ontology-aligned examples improve triple extraction accuracy by up to 44.2%, reduce hallucinations by 22.5%, and enhance format conformance by up to 20.9% compared to baseline LLMs. These findings provide practical insights for prompt construction.
2 Related Work
@@ 3 Benchmark Dataset Creation
We illustrate the process of creating the annotated humanitarian dataset in Fig. 2. We curated 120 online available technical reports: 60 reports from the Cambodian Mine Action Center (CMAC) [^2] https://cmac.gov.kh/publications/ and 60 from Geneva International Centre for Humanitarian Demining (GICHD) websites [^3] https://www.gichd.org/publications-resources/publications/ 1. The reports are all in the PDF format. The curation is based on relevance to the humanitarian demining, language (English), and recency. From the initial humanitarian dataset of 120 PDF documents 1, we filtered a dataset of five m most recent mine action reports from CMAC of overall 233 pages (see Table 1) 2.
Table 1: Dataset Statistics for TextMineX Corpus
@@
Simultaneously, we selected domain-specific ontologies through working with domain experts, by first incorporating data models from the Information Management System for Mine Action (IMSMA) 3. We performed a survey of demining related ontologies together with domain experts and we incorporate six ontologies from Information Management System for Mine Action (IMSMA) Core [^4] https://www.gichd.org/our-response/information-management/imsma-core/. These ontologies are used for demining information system but not strictly for HMA. Furthermore, we add a more general humanitarian domain ontology from Empathi (Gaur et al., 2019) to make the overall HMA ontology more comprehensive. We filter out the concepts that are not relevant to HMA from Empathi by utilizing knowledge of subject-matter expert. For instance, certain concepts related to natural disasters (e.g., flood response) are not directly related to demining. As a result, HMA ontology integrates seven ontologies (160 entity types, 86 relation types) covering diverse aspects of HMA (see details in Sec. 5.1).
The annotation prompt generation 4 first parses reports into text chunks, then systematically combines each chunk with ontology templates. Each template specifies the relation types for a specific domain (e.g., mine action events, land contamination, mine clearance). This generates 2,520 prompts (360 chunks $\times$ 7 templates), where each prompt contains both report context and ontology specifications for that domain.
We randomly select 100 prompts 5 out of the 2,520, and apply them with the same prompts using GPT-4o and Llama3-70B 6 to initially annotate the data as the “LLM annotations 7.
For each prompt and data chunk pair, each LLM generates a set of knowledge triples as outputs. Then, a human annotator who is an expert in natural language processing working with knowledge graphs reviewed these outputs, filtering out incorrect knowledge triples and aggregating the valid ones in a new set. Incorrect knowledge triples can include triples following wrong format, mixed order of entities, hallucinated entities/relations (e.g., triples that are not included in the data chunk) and so on. In addition, there exist duplicate triples which are removed to create the final reference set. Those included 1,095 unique triples across the 100 prompts 8. The end result of the human annotation is a set of knowledge triples which are relevant and correct without hallucination or formatting problems, faithful to the given data chunk. This clean set comprises the “annotated test data” 9. The annotated test data can serve as the ground-truth for evaluations of performance of different techniques and LLMs.
@@ 4.1 Humanitarian Mine Action Task
Figure 3: TextMineX Overview. Reports are preprocessed into paragraph chunks, used as test inputs during inference. Each chunk is combined with an instruction template and ontology, then passed through LLMs for triple extraction. We apply a multi-perspective evaluation using both reference-based and reference-free methods. Extracted triples are stored in a database, queried by developers and HMA domain experts.
@@ 4.3 Multi-Perspective Evaluation
Triple extraction accuracy serves as the primary metric, as it directly measures how well models extract knowledge triples from text. However, accuracy alone does not fully capture model reliability. The hallucination rate evaluates faithfulness by detecting extraneous or fabricated information, while format conformance ensures that outputs adhere to a syntactically valid structure, enabling seamless integration into downstream applications. We employ N-gram matching-based metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) to assess the extracted triples. Additionally, we incorporate BERTScore (Zhang et al., 2019), which leverages word embeddings to capture semantic similarity beyond lexical overlap. To enhance the accuracy of our metrics, we first apply stemming and lemmatization by NLTK to normalize morphological variations. We then compute the accuracy metrics against the manually annotated test set.
@@
We try to find the optimal LLM Judge setting by conducting systematic ranking experiments and analyzing correlations between the LLMs judged ranking and reference-based rankings. For these ranking experiments, we design Judge Prompts that instruct LLMs on evaluation criteria. We use five models: Mistral-7B Jiang et al. (2023), Llama3-8B Grattafiori et al. (2024), Gemma2-9B Team et al. (2024), LLaMA3-70B and GPT-4o Hurst et al. (2024), as extraction models. We rank five responses from five models using GPT-4o, Llama3.1-70B, and Llama3.3-70B as our judge models. The the correlation between different judge models and methods are included in Table 2. The judge prompts follow a fixed template with seven placeholders, where the ontology placeholder represents entity and relation types. Formally, let the input set for the LLM judge prompts be $\{O,C,R_{m_{1}},R_{m_{2}},R_{m_{3}},R_{m_{4}},R_{m_{5}}\}$, where $O$ is the ontology set, $C$ is the set of test contexts, and $R_{m_{1}},\cdots,R_{m_{5}}$ are the five sets of candidate answers from the five extractor models. The judge LLM produces a verdict (output as a ranking):
@@ 5 Experimental Results
We assess the effectiveness of our knowledge triple extraction method through reference-based and reference-free evaluations. Reference-based evaluation compares extracted triples against our curated dataset, while reference-free evaluation relies on LLM judges to assess generated triples without reference data. A key aspect of our analysis is to examine the correlation between these two evaluation paradigms to evaluate the reliability of LLM-based judgments. In addition, we investigate how different prompt strategies influence extraction performance between models.
@@
Figure 4: The combined visualization illustrates the impact of model selection and prompt strategy on extraction performance. Abbreviations on x-axis are about One-shot with: RS = Random Sentences; RP = Random Paragraphs; OS = Ontology-Aligned Sentence; OP = Ontology-Aligned Paragraph. The top Combined Score is achieved by Llama3-70B (93.24) closely followed by GPT-4o (93.13), both with OS prompt setting.
Figure 5 further breaks down the performance of five models across four accuracy evaluation metrics: BLEU, ROUGE, METEOR, and BERTScore. OS demonstration prompts consistently result in the best accuracy across all four metrics and all five models, highlighting their effectiveness for the triple extraction task. BLEU scores peak with OS prompts, with GPT-4o achieving the best performance. ROUGE results show a similar trend, with GPT-4o and Llama3-70B excelling in the OS prompt setting. METEOR follows the same pattern as BLEU and ROUGE, reinforcing the advantages of OS prompts. BERTScore, which measures semantic similarity, shows high clustering across models, suggesting minimal differentiation in performance. Overall, OS demonstration prompts consistently enhance extraction accuracy across all models and metrics.
@@
Figure 7: Number of prompts, entity types, and relation types per the utiized ontologies from IMSMA Core and Empathi.
5.1 Additional Information and Insights for the HMA Ontology
@@ 6 Conclusion
TextMineX addresses the need for knowledge extraction from HMA reports by LLMs and domain ontologies to transform unstructured technical reports into structured knowledge triples. This paper introduces the curated data of HMA reports, ontology, and LLM pipeline, aiming to enable innovations for mine action and natural language processing.
@@
While TextMineX provides the first dataset, evaluation framework and pipeline that demonstrates the feasibility of an LLM-driven extraction pipeline in the specialized domain of humanitarian demining, several limitations are considered to be addressed in the future, including further extension of the dataset. Our evaluation relies on 100 prompt–response examples (yielding 1,095 unique triples).
Assembling and annotating demining data demand extensive domain expertise, so even this modest set offers valuable proof-of-concept insights, but we plan to extend to larger, multilingual collections in future work. Second, this work focuses on reports that are written only in English. In the future, we plan to extend the work to include other languages, because mine-action data are often produced in multiple languages, such as Arabic, French or Ukrainian, depending on the conflict region. Making TextMineX multilingual is, therefore, needed to ensure more inclusiveness.
In this paper, we focus on humanitarian demining and actively collaborate with demining agencies. Thus, the scope of the evaluation framework and pipeline is limited to the demining domain, whereas the applicability of the pipeline in similar domains (e.g, natural disaster management) could be tested as a potential future work.
@@
Humanitarian demining is a high importance decision making domain where incorrect or hallucinated triples can misinform planning and lead to wasted time and resources. To address this, TextMineX combines reference based validation with a bias aware LLM as Judge framework and publishes all prompts and decoding settings for full transparency. We also engage landmine clearance experts throughout development to review and validate outputs. TextMineX is not used to locate mines, as safety remains governed by established GICHD standards [^6] GICHD Mine Action Standards, https://www.gichd.org/our-response/mine-action-standards/, and instead supports expert analysis and planning. By documenting our methods and keeping an expert in the loop, we aim to minimize misinformation and ensure responsible AI deployment in demining operations.
The dataset used in this study consists of publicly available or institutionally provided humanitarian demining reports. These reports were reviewed to ensure they do not contain personally identifiable information (PII) or offensive content. Our usage of the data adheres to privacy standards and is strictly confined to research contexts.
All datasets used in this study were accessed under conditions permitting research use. The curated demining report dataset we constructed is intended solely for academic and research purposes and complies with the original access and licensing conditions. The ontology and pipeline components developed in TextMineX are likewise designed for research and evaluation within humanitarian domains. We do not support or promote deployment of these artifacts in operational or commercial contexts without further validation and ethical review.
AI usage: AI assistance tool is utilized for grammar/spelling check of the paper.